Mary is twice older than her sister.in 5years time she will be 2years older than her sister find how old are they both
Answer by addingup(3677)

Let Mary be x and her sister y. Mary is two times her sister, so:
x = 2y We'll use this value for x next.
In 5 years:
x+5 = (y+5)+2 Substitute for x:
2y+5 = (y+5)+2
2y+3 = y+5
y = 2
------------
Check:
the sister is 2 and Mary is 2*2=4
In 5 years:
2+5=7
4+5=9
and 9-7 = 2, so in 5 years Mary will be 2 years older than sister. CORRECT
:
John